# Evaluating the Robustness of Adversarial Defenses in Malware Detection Systems

### v1

@@ IV-A5 Evaluation Metrics
To rigorously assess the performance and robustness of defenses, we utilize a comprehensive suite of evaluation metrics. These metrics capture both the classification accuracy and resilience to adversarial attacks, and are defined as follows:
@@
  • Attack Success Rate ($\text{ASR}_{k}$): The fraction of adversarial attacks that successfully evade detection, where the Hamming distance between the original and adversarial samples satisfies $d_{H}(\mathbf{x},\mathbf{x}^{\star})\leq k$. This metric measures the model’s resistance to perturbations of varying magnitudes.
@@
IV-A6 Experimental Setup and Configuration
@@
To evaluate the effectiveness of gradient-based attacks, we generate adversarial examples using all 3,104 malicious samples from the test dataset. Each attack is executed with 10,000 iterations of gradient descent, ensuring a thorough exploration of the attack surface. While convergence often occurs earlier, this approach allows for a comprehensive evaluation  [28]. For the CW attack, 8 iterations of binary search are conducted to optimize the penalty weight , maximizing the impact of adversarial perturbations.
As gradient-based attacks produce continuous-valued outputs, the Prioritized Binary Rounding method is employed to enforce binary feature constraints. This post-processing step ensures the generated adversarial examples adhere to the constraints of the domain while maintaining their effectiveness. Hyperparameters for each attack are meticulously tuned for alignment with the defense models under evaluation, ensuring fairness and consistency in the experimental results.
@@
For defenses equipped with adversary detectors, the process is adjusted to use 4 iterations of binary search to identify the smallest effective penalty factor $c$. Additionally, thresholds for identifying suspicious samples are computed by excluding the top 5% of validation examples with the highest confidence, as recommended by [23] and [34]. This strategy ensures that the thresholds are both practical and minimally disruptive to benign classification performance.
IV-B RQ1: Effectiveness of Prioritized Binary Rounding
Figure 1: Robustness comparison of defense models under various binary rounding methods (ASR vs. perturbation budget $k$).
This section evaluates the effectiveness of the proposed Prioritized Binary Rounding method by comparing it with two widely used techniques: Deterministic Binary Rounding and Randomized Binary Rounding. The analysis examines adversarial examples crafted by two distinct and representative attacks: the CW attack, which is an optimization-based method focused on minimizing the $\ell_{2}$-norm of perturbations, and the $\sigma$-zero attack, a sparsity-driven approach specifically targeting the $\ell_{0}$-norm to induce minimal feature modifications. Since the outputs of these attacks are continuous-valued, binary rounding is essential to satisfy the binary feature constraints inherent to binary-constrained domains.
Four defense models were evaluated to ensure comprehensive analysis. The DNN serves as a baseline model without specialized defenses, while AT-rFGSMk, AT-MaxMA, and PAD-SMA (in conservative mode) represent robust defenses known for their effectiveness against adversarial attacks. These models span a spectrum of robustness levels, as further discussed in IV-F.
To ensure a clear comparison, binary-rounded adversarial examples were evaluated alongside their original continuous outputs. The $\ell_{2}$-norm perturbation (characterizing the CW attack) and $\ell_{0}$-norm perturbation (representing the $\sigma$-zero attack) were employed to evaluate the effectiveness of various rounding methods. Figure 1 presents a comparative analysis of the performance of each rounding technique across the defense models.
As depicted in Figure 1, Prioritized Binary Rounding significantly outperforms Deterministic and Randomized Binary Rounding in the $\sigma$-zero attack scenario. It achieves results closely aligned with the original continuous outputs, particularly on the DNN model, where its performance nearly overlaps with the continuous baseline. This effectiveness stems from its alignment with the $\ell_{0}$-norm, which the $\sigma$-zero attack optimizes, and its ability to preserve the adversarial success rate by minimizing the number of feature modifications. Conversely, Randomized Binary Rounding demonstrates the weakest performance, failing to maintain the effectiveness of adversarial examples generated by the $\sigma$-zero attack.
In the CW attack, a more noticeable gap exists between the original continuous outputs and all binary rounding methods. This is attributable to the CW attack’s focus on minimizing the $\ell_{2}$-norm, which often results in numerous small fractional changes, complicating binary conversion. Despite this challenge, Prioritized Binary Rounding still surpasses Deterministic and Randomized Binary Rounding, demonstrating superior robustness across both $\ell_{0}$-norm and $\ell_{2}$-norm-based attacks.
To answer RQ1, Prioritized Binary Rounding outperforms Deterministic and Randomized Binary Rounding by closely preserving the success rates of adversarial attacks across different distance metrics. Its versatility allows for the effective conversion of a broad range of continuous adversarial examples into binary formats, making it a robust and general-purpose solution for binary-constrained domains.
Figure 2: Robustness comparison of defense models under various attack methods (ASR vs. perturbation budget $k$).
IV-C RQ2: Effectiveness of $\sigma$-binary attack
This section evaluates the effectiveness of our proposed $\sigma$-binary attack in comparison with several adversarial attack methods, including the CW attack ($\ell_{2}$-norm-based), the $\sigma$-zero attack ($\ell_{0}$-norm-based), and the Mimicry attack (a gradient-free approach). Additionally, we examine the performance of PGD-$\ell_{1}$, PGD-$\ell_{2}$, and $\ell_{\infty}$, which represent attacks that do not explicitly minimize perturbation. For simplicity, throughout this subsection, PAD-SMA refers exclusively to its configuration in conservative mode.
The results in Table I and Figure 2 demonstrate the superior performance of the $\sigma$-binary attack across all evaluated defenses. Against the baseline DNN, all attacks—except Mimicry—, including $\sigma$-binary, achieve $\text{ASR}_{\infty}$ of 100%, highlighting the vulnerability of DNNs to adversarial perturbations. However, significant differences arise when robust defenses such as AT-rFGSMk, AT-MaxMA, and PAD-SMA are evaluated.
For AT-rFGSMk, $\sigma$-binary achieves an impressive $\text{ASR}_{\infty}$ of 99.45%, followed by PGD-$\ell_{2}$ (89.21%) and $\sigma$-zero (85.95 %). On AT-MaxMA, $\sigma$-binary reaches $\text{ASR}_{\infty}$ of 96.62%, outperforming $\sigma$-zero (92.56%) and PGD-$\ell_{2}$ (91.75%). Against the strongest defense, PAD-SMA, $\sigma$-binary achieves an $\text{ASR}_{\infty}$ of 94.56%, significantly surpassing PGD-$\ell_{2}$ (79.90%) and $\sigma$-zero (79.64%). In contrast, PGD-$\ell_{1}$ struggles to adapt to robust defenses, with a low $\text{ASR}_{\infty}$ of 32.76% on AT-MaxMA, indicating limited versatility.
The CW attack’s poor performance after binarization is noteworthy. Designed to minimize $\ell_{2}$-norm perturbations, CW produces adversarial examples near the decision boundary, making it particularly vulnerable to binarization, which disrupts its finely tuned perturbations. In contrast, PGD-$\ell_{2}$ and $\ell_{\infty}$ avoid this limitation by producing perturbations that push further beyond the decision boundary, resulting in higher success rates post-binarization.
Table I: Accuracy (%) of defenses under adaptive attacks.
Attack name $\text{ASR}_{\infty}$ (%)
DNN AT-rFGSMk AT-MaxMA PAD-SMA
CW (binarized) 100.00 68.36 65.37 36.82
$\sigma$-zero (binarized) 100.00 85.95 92.56 79.64
Mimicry 99.81 83.92 83.09 67.20
PGD-$\ell_{1}$ 100.00 79.96 32.76 69.14
PGD-$\ell_{2}$ (binarized) 100.00 89.21 91.75 79.90
$\ell_{\infty}$ (binarized) 100.00 85.47 84.09 75.52
$\sigma$-binary 100.00 99.45 96.62 94.56
As shown in Figure 2, $\sigma$-binary consistently achieves higher ASRs at smaller perturbation budgets compared to other attacks, demonstrating its efficiency in generating effective adversarial examples with minimal input modifications. Although PGD-$\ell_{2}$ and $\ell_{\infty}$ improve slightly at larger perturbation budgets, they remain less effective than $\sigma$-binary, especially against PAD-SMA. Mimicry, a gradient-free attack, performs poorly against PAD-SMA due to the robust adversarial detection capabilities of PAD-SMA, particularly against high-perturbation attacks.
Answer to RQ2: The $\sigma$-binary attack is highly effective across a wide range of defenses, including PAD-SMA. It consistently achieves higher ASRs with smaller perturbation budgets, outperforming all other evaluated attacks and establishing itself as a robust and efficient adversarial attack method.
IV-D RQ3: Baseline performance of defenses

### v2

@@ V-A5 Evaluation Metrics
To rigorously assess the performance and robustness of defenses, we utilize a comprehensive suite of evaluation metrics. These metrics capture both the classification accuracy and resilience to adversarial attacks, and are defined as follows.
@@
  • Attack Success Rate ($\text{ASR}_{k}$): The proportion of adversarial attacks that successfully evade detection, where the Hamming distance between the original and adversarial samples satisfies $d_{H}(\mathbf{x},\mathbf{x}^{\star})\leq k$. This metric measures the model’s resistance to perturbations of varying magnitudes.
@@
  • Interquartile Range (IQR): The dispersion of Hamming distances among successful attacks, reflecting the stability and consistency of perturbation magnitudes.
@@
V-B Experimental setup
V-B1 Hardware, software and reproducibility
Experiments were executed on a workstation running Microsoft Windows 11 Pro with an 11th Gen Intel Core i5-11500H CPU (6 cores, 12 threads), 32 GB RAM, and an NVIDIA RTX A2000 Laptop GPU (4 GB VRAM). The experimental codebase uses Python 3.12.5 and PyTorch 2.5.0 with CUDA 12.4. To facilitate reproducibility, all software dependencies, hyperparameters, and processed datasets (including the exact train/validation/test splits) are publicly available.
V-B2 Data splits and protocol
All experiments use the MalScan dataset with fixed partitions: 60% training, 20% validation, and 20% test. Identical splits and model initializations were maintained across experiments to ensure fair comparisons. Performance is reported on the full test set, which remained strictly held-out during training and validation.
V-B3 Model selection and hyperparameters
Where available, defense models adopt the hyperparameter settings reported in their original works. This choice avoids confounding differences due to re-tuning and enables a direct assessment of comparative robustness under the conditions validated by prior studies.
V-B4 Numerical stability and bias control
To limit numerical instability, gradient clipping was applied during training. Gradient-based attacks use normalized step sizes to ensure consistent perturbation scaling across defenses. Experimental bias was further reduced by fixing random seeds for Python, NumPy, and PyTorch.
V-B5 Attack configuration
@@
To evaluate the effectiveness of gradient-based attacks, we generate adversarial examples using all 3,104 malicious samples in the test dataset. For attack-strength comparisons, we allowed up to 1,000 gradient-descent iterations; for robustness evaluation we extended the budget to 10,000 iterations of gradient descent, ensuring a thorough exploration of the attack surface. While convergence often occurs earlier, this approach allows for a comprehensive evaluation [carlini2017towards]. For the CW attack, 8 iterations of binary search are conducted to optimize the penalty weight c, maximizing the impact of adversarial perturbations. Continuous outputs from gradient-based attacks were converted to the binary domain via deterministic rounding (threshold $0.5$), except where attacker formulations natively operate over binary variables. For the $\sigma$-binary attack we apply the PBR scheme to preserve adversarial efficacy while enforcing binary constraints.
V-B6 Detector calibration
@@
For defenses equipped with adversary detectors, the process is adjusted to use 4 iterations of binary search to identify the smallest effective penalty factor $c$. Additionally, thresholds for identifying suspicious samples are computed by excluding the top 5% of validation examples with the highest confidence, as recommended by [li2023pad] and [sperl2020dla]. This strategy ensures that the thresholds are both practical and minimally disruptive to benign classification performance.
The above configuration establishes a standardized and reproducible evaluation pipeline for all experiments. The following section presents empirical results addressing RQ1–RQ5, highlighting the comparative robustness and efficiency of the proposed $\sigma$-binary framework.
V-C RQ1: Performance of $\sigma$-binary attack
Table II: Attack Success Rate (ASR, %) under four defense models. Columns report ASR for 10, 20, 50, 100, and $\infty$ perturbation budgets, along with mean runtime per sample $\bar{t}$ (in seconds/sample) and maximum VRAM usage (in GB). The best result per row is in bold.
Attack DNN AT-rFGSMk
ASR10 ASR20 ASR50 ASR100 ASR $\bar{t}$ VRAM ASR10 ASR20 ASR50 ASR100 ASR $\bar{t}$ VRAM
BCA 93.2 99.7 99.8 99.8 99.8 0.008 0.6 29.5 48.9 67.4 73.3 75.2 0.011 0.9
BGA 2.1 2.1 2.1 2.1 100.0 0.001 0.8 1.6 4.8 16.9 37.5 50.0 0.015 1.0
Grosse 93.1 99.7 99.8 99.8 99.8 0.009 0.6 29.5 49.1 67.5 73.7 75.6 0.012 0.9
rFGSM 2.1 2.1 2.1 26.0 99.7 0.008 0.7 1.1 1.2 3.4 14.3 79.5 0.011 0.9
PGD-$\ell_{1}$ 93.2 99.7 99.8 99.8 99.8 0.012 0.7 39.1 57.3 73.8 78.3 79.5 0.018 0.9
PGD-$\ell_{2}$ 2.1 2.1 2.1 2.1 93.8 0.012 0.7 1.1 1.2 4.1 21.2 77.7 0.012 0.9
PGD-$\ell_{\infty}$ 2.1 2.1 2.1 2.1 99.8 0.012 0.7 1.1 1.2 4.4 20.6 84.5 0.012 0.9
iMaxMA 2.1 2.1 2.1 2.1 99.8 0.036 0.8 1.1 1.1 4.0 20.1 88.8 0.045 1.0
SMA 93.2 99.7 99.8 99.8 99.8 0.011 0.8 39.2 57.7 70.9 72.3 72.4 0.025 1.1
Mimicry 16.6 35.1 77.8 98.4 99.8 0.012 1.0 12.2 23.1 47.3 62.3 73.6 0.011 1.2
$\sigma$-zero 89.4 94.7 94.9 94.9 94.9 0.028 0.8 42.8 56.4 70.0 73.7 73.8 0.027 0.9
CW 43.7 51.4 51.6 51.6 51.6 0.107 1.9 20.8 29.9 38.7 43.2 44.6 0.105 2.2
$\sigma$-binary 94.5 99.8 100.0 100.0 100.0 0.046 1.0 48.2 67.0 90.3 98.8 99.4 0.067 1.2
Attack AT-MaxMA PAD-SMA
ASR10 ASR20 ASR50 ASR100 ASR $\bar{t}$ VRAM ASR10 ASR20 ASR50 ASR100 ASR $\bar{t}$ VRAM
BCA 4.9 6.2 9.6 15.3 30.5 0.011 0.9 2.6 4.2 13.2 15.8 16.3 0.074 0.9
BGA 1.3 1.8 2.8 3.0 3.5 0.014 1.1 1.5 1.5 1.5 1.5 1.5 0.088 1.2
Grosse 4.9 6.2 9.5 15.0 28.8 0.012 0.9 2.9 4.3 13.4 16.4 17.0 0.076 0.9
rFGSM 0.8 0.9 2.5 9.2 62.9 0.011 1.0 0.8 1.0 2.3 6.6 11.8 0.074 1.0
PGD-$\ell_{1}$ 5.6 7.7 12.4 18.8 30.8 0.018 1.0 2.3 5.2 14.4 18.7 20.6 0.101 1.0
PGD-$\ell_{2}$ 0.8 0.8 3.8 12.7 69.2 0.012 0.9 0.8 1.0 2.3 9.3 16.2 0.079 1.0
PGD-$\ell_{\infty}$ 0.8 0.9 3.3 14.8 73.9 0.012 0.9 0.8 0.9 2.6 9.4 15.3 0.076 1.0
iMaxMA 0.8 0.9 2.6 11.8 86.9 0.052 1.1 1.0 1.7 4.1 13.0 22.4 0.439 1.1
SMA 9.2 11.4 17.6 29.9 35.4 0.044 1.1 3.8 5.7 13.9 14.7 14.7 0.340 1.1
Mimicry 8.8 18.2 37.8 57.5 66.1 0.011 1.2 11.1 18.9 47.9 57.4 59.5 0.018 1.3
$\sigma$-zero 10.2 14.5 30.0 42.2 45.0 0.027 1.0 25.8 49.2 65.4 68.2 68.3 0.151 1.2
CW 7.7 12.4 22.9 34.1 41.6 0.105 2.2 5.9 11.5 22.5 24.4 24.4 0.696 2.3
$\sigma$-binary 21.9 30.5 62.5 85.7 96.6 0.068 1.1 29.2 57.4 84.5 92.2 93.9 0.448 1.3
Table II presents a comparative analysis of the proposed $\sigma$-binary attack against a wide range of state-of-the-art adversarial methods under four representative defense models: DNN, AT-rFGSMk, AT-MaxMA, and PAD-SMA. These models were selected as they represent the most robust and widely adopted defenses on the MalScan dataset—namely, AT-rFGSMk, AT-MaxMA, and PAD-SMA—alongside the standard DNN model as a baseline. Complete results for additional defense mechanisms are reported in Appendix A-B.
The table reports the Attack Success Rate ($\text{ASR}_{k}$) under multiple feature-modification budgets $k\in\{10,20,50,100,\infty\}$, where each $k$ defines the Hamming distance limit $d_{H}(\mathbf{x},\mathbf{x}^{\star})$ between original and adversarial samples. We also include mean runtime ($t_{\text{mean}}$) and maximum GPU memory usage (VRAM) across all 3,104 malware samples to assess computational efficiency. All experiments follow the standardized setup in Section V-B using identical hardware, software, and batch configurations to ensure fair and reproducible comparisons.
Effectiveness
Consistent superiority across defenses.
The proposed $\sigma$-binary attack achieves the highest $\text{ASR}_{k}$ across all four defenses and modification budgets. On the baseline DNN, it reaches near-perfect evasion even under tight budgets ($\text{ASR}_{10}=94.5\%$, saturating at $100\%$ for $k\geq 50$), outperforming strong baselines such as PGD-$\ell_{1}$ and SMA. Under adversarially trained defenses, $\sigma$-binary maintains its advantage—achieving $\text{ASR}_{50}=90.3\%$ and $\text{ASR}_{100}=98.8\%$ against AT-rFGSMk, and $\text{ASR}_{100}=85.7\%$ under AT-MaxMA—while sustaining robust performance ($\text{ASR}_{100}=92.2\%$) even against the detector-based PAD-SMA.
Robustness to binarization.
Attacks originally optimized in continuous space (e.g., CW, $\sigma$-zero) show sharp degradation after rounding, as their fine-grained perturbations near decision boundaries are disrupted by binary projection. In contrast, $\sigma$-binary’s discrete-aware optimization and prioritized rounding maintain stability across defenses, producing consistently higher ASR values and confirming the importance of binary-native optimization for malware domains.
Strengthened Mimicry remains non-competitive.
[…]
Stable performance under tight budgets.
While competing methods show inconsistent trends across defenses and budgets, $\sigma$-binary maintains stable high performance even with minimal feature modifications ($k=10$), highlighting the efficiency of its distance-aware optimization and rounding strategy.
Figure 1: Attack Success Rate (ASR) versus perturbation budget $k$ (Hamming distance $d_{H}$) across four defense models. Only the most competitive attacks are shown for clarity.
Reliability and Convergence Behavior
Complementary to Table II, we present the robustness evaluation curves in Figure 1 for each attack under four representative defenses. Similar curves for all remaining defense models are included in Appendix A-B. Due to the large number of evaluated attacks, only the most competitive methods are shown in Figure 1 for clarity, ensuring readability while preserving the key comparative trends across defenses. Across all defenses, $\sigma$-binary exhibits a steep and stable growth curve, saturating near $100\%$ for $k<100$, while alternatives such as SMA and Mimicry show slower or irregular improvements. Under the strongest defenses (PAD-SMA and AT-MaxMA), $\sigma$-binary consistently outperforms competitors by 20–30 percentage points, demonstrating stable convergence and high reliability across perturbation ranges. Unlike gradient-based baselines that often saturate early or fail to converge fully, $\sigma$-binary almost always finds valid adversarial solutions, ensuring faithful robustness evaluation.
Efficiency
Despite its stronger optimization framework, $\sigma$-binary remains computationally practical. Its mean runtime is moderately higher than lightweight attacks due to nested optimization and adaptive rounding but remains computationally efficient overall. For instance, under PAD-SMA, the proposed $\sigma$-binary attack completes inference with a mean runtime of approximately $0.45\penalty 10000\ \text{s/sample}$ while using only $1.3\penalty 10000\ \text{GB}$ of VRAM—demonstrating practical computational efficiency even on a standard mobile workstation. This balance between effectiveness and efficiency demonstrates that $\sigma$-binary’s robustness does not come at the cost of prohibitive resource requirements.
Ablation Study on $\sigma$-binary
To further investigate the contribution of each component in the proposed $\sigma$-binary framework, we conducted an ablation study under two representative adversarially trained defenses: AT-rFGSMk and AT-MaxMA. Each variant disables one major component of the framework while keeping all other parameters fixed, allowing an isolated assessment of its impact on attack performance. The examined components include: (i) the differentiable Hamming distance (${d}_{H}$) approximation, (ii) the gradient normalization factor, (iii) the adaptive sparsity projection operator, (iv) the Prioritized Binary Rounding (PBR) mechanism, and (v) the dynamic confidence adjustment strategy. Performance is measured using attack success rates $\text{ASR}_{k}$ at modification budgets $k\in\{10,20,50,100,\infty\}$, as well as the median and interquartile range (IQR) of perturbation magnitudes $\tilde{d}_{H}$ among successful attacks. To ensure stability and comparability, all experiments in this ablation were performed on a fixed subset of 1,000 randomly sampled test instances with an extended optimization budget of 10,000 iterations. The detailed results are reported in Table III.
Table III: Ablation study of the proposed $\sigma$-binary attack under two adversarially trained defenses (AT-rFGSMk and AT-MaxMA). Each variant removes one core component while keeping all others unchanged. $\text{ASR}_{k}$ denotes the attack success rate (%) under a Hamming distance constraint $d_{H}(\mathbf{x},\mathbf{x}^{\star})\leq k$. $\tilde{d}_{H,\text{median}}$ and IQR represent the median and interquartile range of perturbation magnitudes among successful attacks. Bold values indicate the best performance within each defense.
Configuration AT-rFGSMk AT-MaxMA
ASR10 ASR20 ASR50 ASR100 ASR Med / IQR ASR10 ASR20 ASR50 ASR100 ASR Med / IQR
Full $\sigma$-binary (all) 51.1 69.0 91.9 98.9 99.4 10.0 / 19.0 23.8 36.9 64.3 86.9 97.0 31.0 / 53.0
Without ${d}_{H}$ approximation 7.5 19.8 40.3 75.0 99.3 68.0 / 73.0 1.8 3.2 12.5 17.5 94.9 215.0 / 116.0
Without normalization 45.0 63.1 87.6 98.2 99.1 12.0 / 26.0 18.7 25.8 59.5 85.6 96.9 42.0 / 52.0
Without adaptive projection 50.4 68.2 91.2 98.9 99.4 10.0 / 21.0 23.1 36.6 64.2 86.7 97.0 33.0 / 55.0
Without PBR rounding 50.4 68.8 91.8 98.9 99.4 10.0 / 20.0 23.0 36.0 63.1 86.7 97.0 32.0 / 54.0
Without dynamic confidence 50.0 68.5 91.9 98.9 99.4 10.0 / 20.0 23.3 34.8 64.6 86.9 97.0 35.0 / 54.5
As shown in Table III, each component contributes meaningfully to the performance of the $\sigma$-binary framework. Removing the differentiable ${d}_{H}$ approximation causes the largest drop in low-budget success rates ($k\leq 50$) and a substantial increase in perturbation distance, confirming its key role in stabilizing gradient-based optimization. Disabling gradient normalization factor or adaptive projection slightly reduces $\text{ASR}_{k}$ and increases variability, indicating that both components help maintain balanced updates and sparsity. When the PBR mechanism is replaced with deterministic binary rounding, the results remain relatively strong—thanks to checkpointing of the best rounded candidates—but perturbations become less sparse and slightly larger. Finally, omitting dynamic confidence adjustment yields comparable final ASR values but broader perturbation dispersion (higher IQR), suggesting less consistent convergence. Overall, the results confirm that the combination of smooth distance modeling, adaptive sparsity, and prioritized rounding is crucial for achieving the high efficiency and reliability of the full $\sigma$-binary attack.
Ablation Study on PBR
To quantify the contribution of the feature-importance weighting mechanism in the proposed PBR framework, we performed an ablation study comparing two variants: (i) a baseline configuration that prioritizes features solely according to perturbation magnitude (derived from gradient magnitude), and (ii) an enhanced configuration that additionally incorporates discriminative feature importance derived from the training data.
In the enhanced version, discriminative feature importance acts as a secondary prioritization criterion: when multiple features share the same gradient magnitude, their order is resolved according to discriminative importance. Consequently, its overall impact remains modest, but it can provide incremental improvements in cases where gradient information alone is insufficient to distinguish between features of equal significance.
Table IV: Ablation study on the effect of incorporating feature importance in PBR. $\text{ASR}_{k}$ denotes the attack success rate (%) under a Hamming distance constraint $d_{H}(\mathbf{x},\mathbf{x}^{\star})\leq k$. $\tilde{d}_{H,\mathrm{median}}$ and IQR represent the median and interquartile range of normalized Hamming distances across successful attacks.
Model Iter. Method $\text{ASR}_{10}$ $\text{ASR}_{20}$ $\text{ASR}_{50}$ $\text{ASR}_{100}$ $\text{ASR}_{\infty}$ $\tilde{d}_{H,\mathrm{median}}$ IQR
AT-rFGSMk 100 Perturbation Magnitude $32.02$ $49.48$ $74.48$ $90.43$ $92.20$ $18.00$ $33.00$
+ Feature Importance 32.15 49.55 74.52 90.46 $92.20$ $18.00$ $33.00$
1000 Perturbation Magnitude $47.94$ $66.95$ $90.30$ $98.81$ $99.39$ $11.00$ $22.00$
+ Feature Importance $47.94$ $66.95$ $90.30$ $98.81$ $99.39$ $11.00$ $22.00$
AT-MaxMA 100 Perturbation Magnitude $15.79$ $22.97$ $53.16$ $84.54$ $95.52$ $47.00$ $55.0$
+ Feature Importance 15.88 23.10 $53.16$ $84.54$ $95.52$ $47.00$ $51.00$
1000 Perturbation Magnitude $21.97$ $30.48$ $62.50$ $85.76$ $96.59$ $38.00$ $54.00$
+ Feature Importance $21.97$ $30.48$ $62.50$ $85.76$ $96.59$ $38.00$ $54.00$
Both variants share identical hyperparameter settings and are evaluated on a fixed subset of 1,000 randomly sampled test instances to ensure fairness. Each experiment was run for 100 and 1,000 optimization iterations to examine both early-stage and converged behavior. Table IV reports the attack success rates ($\text{ASR}_{k}$) under different perturbation budgets $k\in\{10,20,50,100,\infty\}$, together with the median and interquartile range (IQR) of normalized Hamming distances $(\tilde{d}_{H,\mathrm{median}},\mathrm{IQR})$ among successful attacks.
At smaller iteration budgets ($100$), incorporating discriminative feature importance yields consistent though marginal gains of approximately 0.1–0.2 percentage points in $\text{ASR}_{k}$ across most perturbation thresholds while maintaining nearly identical distance distributions. This indicates that secondary ordering by discriminative importance helps guide the rounding process toward semantically meaningful binary features, improving optimization efficiency in some instances.
When the iteration budget increases to 1,0001{,}000, both variants converge to virtually identical performance. The ASR, median distances, and IQR values become indistinguishable, suggesting that with sufficient optimization steps, the attack inherently learns to emphasize salient features even without explicit feature-importance guidance.
Overall, these findings show that incorporating discriminative feature importance as a secondary tie-breaking criterion can slightly improve early-stage optimization but has negligible influence after full convergence, thereby confirming that PBR’s primary effectiveness stems from gradient-based prioritization.
Hyperparameter Robustness of the $\sigma$-Binary Attack
We examined the sensitivity of the proposed $\sigma$-binary attack to its key hyperparameters: the smoothing factor $\sigma$, the initial sparsity threshold $\gamma_{0}$, and the sparsity adjustment rate $t$. All experiments were performed on 1,000 randomly selected test samples for both AT-rFGSMk and AT-MaxMA models.
For the ablation on the sparsity adaptation rate $t$, the other hyperparameters were fixed to their default values of $\gamma_{0}=0.3$ and $\sigma=10^{-3}$. In contrast, in the preceding experiments examining the effects of $\gamma_{0}$ and $\sigma$, these parameters were systematically varied within the ranges $\gamma_{0}\in[0,0.5]$ and $\sigma\in[10^{-6},10^{-1}]$, respectively, while the sparsity adaptation rate was held constant at its default value of $t=0.01$. For each configuration, we report the Attack Success Rate (ASR) at multiple perturbation budgets and the median normalized Hamming distance ($\tilde{d}_{H,\mathrm{median}}$).
Effect of $\gamma_{0}$ and $\sigma$: Fig. 2 summarizes the interaction between the smoothing parameter $\sigma$ and the initial sparsity threshold $\gamma_{0}$. Across all tested values, the attack performance remains remarkably stable. Both AT-rFGSMk and AT-MaxMA exhibit consistent ASR and $\tilde{d}_{H,\mathrm{median}}$, confirming that the adaptive projection operator $\Pi_{\gamma}$ rapidly neutralizes the impact of $\gamma_{0}$. Similarly, moderate values of $\sigma$ (between $10^{-6}$ and $10^{-2}$) yield comparable results. Only very large $\sigma$ values ($10^{-1}$) slightly degrade ASR at small perturbation budgets and increase $\tilde{d}_{H,\mathrm{median}}$, due to excessive smoothing of the Hamming surrogate that reduces sparsity in early optimization steps.
Effect of the sparsity adjustment factor $t$: Fig. 3 illustrates the impact of the sparsity update rate $t$. The $\sigma$-binary attack remains robust for $t\leq 10^{-1}$, maintaining high ASR and low $\tilde{d}_{H,\mathrm{median}}$ across both defenses. For larger $t$, AT-MaxMA becomes more sensitive, showing a drop in ASR and an increase in $\tilde{d}_{H,\mathrm{median}}$. This behavior indicates that overly aggressive updates of $\gamma$ can destabilize the optimization process, leading to less compact perturbations.
Note on dynamic confidence adjustment: We did not include a separate sensitivity analysis for the dynamic confidence adjustment parameters, as this mechanism becomes active only after convergence of the loss function. Its role is to refine already converged adversarial examples by slightly perturbing the confidence margins around the decision boundary, without altering the loss or optimization trajectory. As shown in Appendix III, the effect of this mechanism is marginal—primarily improving results under high iteration budgets and stronger defenses by stabilizing late-stage convergence. Because it does not materially influence the main optimization trajectory, its hyperparameter tuning has negligible practical impact and is therefore excluded from this robustness study.
Discussion: Overall, the $\sigma$-binary attack demonstrates strong robustness to its principal hyperparameters. A broad configuration range produces near-optimal performance, enabling efficient deployment without per-instance tuning. Empirically, we recommend the following default ranges:
$\sigma\in[10^{-6},10^{-2}],\quad t\leq 10^{-1},\quad\gamma_{0}\in[0,0.5].$
These settings consistently yield high ASR with minimal perturbation size for both evaluated defenses.
Figure 2: Hyperparameter robustness of the $\sigma$-binary attack across smoothing parameter $\sigma$ and initial sparsity threshold $\gamma_{0}$. Each cell shows the Attack Success Rate (ASR) at different perturbation budgets and the corresponding median normalized Hamming distance. Top panel: AT-rFGSMk; bottom panel: AT-MaxMA.
Figure 3: Effect of the sparsity adjustment factor $t$ on the $\sigma$-binary attack. Left: ASR versus $t$ (log scale). Right: median normalized Hamming distance $\tilde{d}_{H,\mathrm{median}}$ versus $t$.
Answer to RQ1: $\sigma$-binary achieves consistently higher attack success rates across all defenses while maintaining practical efficiency and remarkable stability. Its advantage stems from the integrated design of differentiable distance modeling, adaptive sparsity control, and prioritized rounding, which together enable precise and reliable optimization in binary feature spaces. The attack remains robust to hyperparameter variation and maintains convergence stability even under strong adversarial training, providing both superior effectiveness and trustworthy evaluation of malware defense robustness.
V-D RQ2: Baseline performance of defenses